# From Transparent to Opaque: Rethinking Neural Implicit Surfaces with $\alpha$-NeuS

### v1

@@ Abstract
Traditional 3D shape reconstruction techniques from multi-view images, such as structure from motion and multi-view stereo, primarily focus on opaque surfaces. Similarly, recent advances in neural radiance fields and its variants also primarily address opaque objects, encountering difficulties with the complex lighting effects caused by transparent materials. This paper introduces $\alpha$-NeuS, a new method for simultaneously reconstructing thin transparent objects and opaque objects based on neural implicit surfaces (NeuS). Our method leverages the observation that transparent surfaces induce local extreme values in the learned distance fields during neural volumetric rendering, contrasting with opaque surfaces that align with zero level sets. Traditional iso-surfacing algorithms such as marching cubes, which rely on fixed iso-values, are ill-suited for this data. We address this by taking the absolute value of the distance field and developing an optimization method that extracts level sets corresponding to both non-negative local minima and zero iso-values. We prove that the reconstructed surfaces are unbiased for both transparent and opaque objects. To validate our approach, we construct a benchmark that includes both real-world and synthetic scenes, demonstrating its practical utility and effectiveness. Our data and code are publicly available at %Number␣of␣real␣datahttps://github.com/728388808/alpha-NeuS.
@@ 1 Introduction
Surface reconstruction from multi-view images has been an important area of research for decades. Traditional methods such as structure from motion (SfM) [1] and multi-view stereo (MVS) [2] calibrate images and reconstruct 3D geometry based on color consistency. Recently, the emergence of Neural Radiance Fields (NeRF) [3] has revolutionized the field, producing impressive results in novel view synthesis results via volume rendering. Its implicit surface-based variants, such as NeuS [4], VolSDF [5], HF-NeuS [6], and NeuS2 [7], further advance this field by reconstructing high-quality geometry and appearance through the learning of signed distance fields (SDFs). However, these methods are limited to reconstructing opaque surfaces. Reconstructing transparent surfaces presents a greater challenge, with relatively few investigations to date.
@@
    We present a method for simultaneously extracting the unbiased surfaces corresponding to the target geometry of transparent objects and opaque objects, from mixed SDF and unsigned distance field (UDF).
@@ 2 Related Works
Reconstruction of transparent objects is much more complicated. Complex light paths [30] caused by refraction and reflection in transparent object prevent multiview stereo from solving this problem [31]. Traditional methods [32, 33, 34] use additional devices or assumptions to reconstruct transparent objects. Li et al. [26] used deep learning to further improve the quality of reconstruction without additional inputs, but tend to produce over-smoothing results. With the development of NeRF [3], some works have explored how to use neural rendering to reconstruct transparent objects [35, 36, 25, 37, 27] for capturing more details. However, they all just focus on transparent objects neglecting opaque objects.
@@
But none of them can reconstruct both transparent and opaque objects. They all have their own assumptions or conditions for reducing the effect of reflection or refraction. It is non-trivial to combine the two tasks. Please refer to Table 1 for a comprehensive comparison on the use cases. A concurrent work [39] proposed a NeRF-based efficient rendering method for non-opaque scenes with baked quadrature fields. Another concurrent work, $\alpha$Surf [29], extends Plenoxels [40] for modeling both transparent objects and opaque objects, while ignoring the effect of reflection and refraction. In this paper, we propose a new algorithm to reconstruct thin transparent objects and opaque objects uniformly based on NeuS [4].
@@ 3.1 Unbiased density mapping in NeuS across opacities
Figure 1: Conceptual illustration of the signed distances along a ray (black horizontal line) through a scene containing a single object (represented by a vertical line segment). (a) When the rendered opacity $\alpha$ is less than $0.5$, both the front and back faces of the object coincide with each other, aligning the maximal weight with the local minimum of the distance field. (b) When $\alpha$ exceeds $0.5$ but less than 1, the back face, which is not rendered, is separated from the front face. The further away the back face is, the more opaque the rendered front face is. The maximum weight in this case is aligned with the position of zero distance values. (c) For a fully opaque surface, the back face is infinitely away. The scene can therefore be considered the single ray-plane intersection discussed by NeuS [4].
@@ 5 Conclusion
This work was supported in part by the Basic Research Project of ISCAS (ISCAS-JCMS-202303), in part by the Major Research Project of ISCAS (ISCAS-ZD-202401), in part by the National Key R&D Program of China (2023YFB3002901), in part by the Ministry of Education, Singapore, under its Academic Research Fund Grants (MOE-T2EP20220-0005 & RT19/22) and the RIE2020 Industry Alignment Fund–Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).

### v2

@@ Abstract
Traditional 3D shape reconstruction techniques from multi-view images, such as structure from motion and multi-view stereo, face challenges in reconstructing transparent objects. Recent advances in neural radiance fields and its variants primarily address opaque or transparent objects, encountering difficulties to reconstruct both transparent and opaque objects simultaneously. This paper introduces $\alpha$-NeuS—an extension of NeuS—that proves NeuS is unbiased for materials from fully transparent to fully opaque. We find that transparent and opaque surfaces align with the non-negative local minima and the zero iso-surface, respectively, in the learned distance field of NeuS. Traditional iso-surfacing extraction algorithms, such as marching cubes, which rely on fixed iso-values, are ill-suited for such data. We develop a method to extract the transparent and opaque surface simultaneously based on DCUDF. To validate our approach, we construct a benchmark that includes both real-world and synthetic scenes, demonstrating its practical utility and effectiveness. Our data and code are publicly available at %Number␣of␣real␣datahttps://github.com/728388808/alpha-NeuS.
@@ 1 Introduction
Surface reconstruction from multi-view images has been an important area of research for decades. Traditional methods such as structure from motion (SfM) [1] and multi-view stereo (MVS) [2] calibrate images and reconstruct 3D geometry based on color consistency. Recently, the emergence of Neural Radiance Fields (NeRF) [3] has revolutionized the field, producing impressive results in novel view synthesis results via volume rendering. Its implicit surface-based variants, such as NeuS [4], VolSDF [5], HF-NeuS [6], and NeuS2 [7], further advance this field by reconstructing high-quality geometry and appearance through the learning of signed distance fields (SDFs). However, these NeRF family methods are limited to reconstructing opaque surfaces. Reconstructing transparent surfaces presents a greater challenge, with relatively few investigations to date.
@@
    We present a method for simultaneously extracting the unbiased surfaces corresponding to the target geometry of transparent objects and opaque objects based on DCUDF [11], from mixed SDF and unsigned distance field (UDF).
@@ 2 Related Works
Reconstruction of transparent objects presents a significant challenge due to the complex light paths [30] caused by refraction and reflection, which hinder multiview stereo from effectively solving this problem [31]. Traditional methods [32, 33, 34] use additional devices or assumptions to reconstruct transparent objects. Li et al. [26] used deep learning to further improve the quality of reconstruction without additional inputs, but tend to produce over-smoothing results. With the development of NeRF [3], some works have explored how to use neural rendering to reconstruct transparent objects [35, 36, 25, 37, 27] for capturing more details. However, they all just focus on transparent objects neglecting opaque objects.
@@
But none of them can reconstruct both transparent and opaque objects. They all have their own assumptions or conditions for reducing the effect of reflection or refraction. It is non-trivial to combine the two tasks. Please refer to Table 1 for a comprehensive comparison on the use cases. A concurrent work [39] proposed a NeRF-based efficient rendering method for non-opaque scenes with baked quadrature fields. Another concurrent work, $\alpha$Surf [29], extends Plenoxels [40] for modeling both transparent objects and opaque objects, while ignoring the effect of reflection and refraction. Concurrently, the work NU-NeRF [41] tries to model both transparent and opaque objects while recovering refractions. In this paper, we propose a new algorithm to reconstruct thin transparent objects and opaque objects uniformly based on NeuS [4].
@@ 3.1 Unbiased density mapping in NeuS across opacities
Figure 1: Conceptual illustration of the signed distances along a ray (black horizontal line) through a scene containing a single object (represented by a vertical line segment). (a) When the rendered opacity $\alpha$ is less than or equal to $0.5$, both the front and back faces of the object coincide with each other, aligning the maximal weight with the local minimum of the distance field. (b) When $\alpha$ exceeds $0.5$ but less than 1, the back face, which is not rendered, is separated from the front face. The further away the back face is, the more opaque the rendered front face is. The maximum weight in this case is aligned with the position of zero distance values. (c) For a fully opaque surface, the back face is infinitely away. The scene can therefore be considered the single ray-plane intersection discussed by NeuS [4].
@@ 5 Conclusion
This work was supported in part by the National Key R&D Program of China (2023YFB3002901), in part by the Basic Research Project of ISCAS (ISCAS-JCMS-202303), in part by the Major Research Project of ISCAS (ISCAS-ZD-202401), in part by the Ministry of Education, Singapore, under its Academic Research Fund Grants (MOE-T2EP20220-0005 & RT19/22) and the RIE2020 Industry Alignment Fund–Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).